OpenStudy (anonymous):

Describe each of the following set by listing the elements in the set: {x ∈ ℚ | x(x2 – 4)(2x + 3) = 0}

OpenStudy (anonymous):

@mathmate hey can you work through this with me? the answer is supposed to be {-2, -3/2, 0, 2} and i got 4 :(

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

@freckles

OpenStudy (mathmate):

We need to use the zero product property (see for example https://www.mathsisfun.com/algebra/zero-product-property.html ) If (x^2-4)x=0, which can be factored as (x+2)(x-2)x=0. Using the zero product property, then (x+2)=0, or (x-2)=0, or x=0 (note: the or's above are inclusive or's, meaning and/or ). So if (x+2)=0, then x=-2, if (x-2)=0, then x=2 if x=0, then x=0. Apply the zero product property to your problem and you'll get the correct answers.